x=(1/4)y , y=(1/5)z so what is equal of x:y:z?
Answer by ramkikk66(644)


x=(1/4)y , y=(1/5)z so what is equal of x:y:z?
x = (1/4)y. i.e x/y = 1/4 or x:y = 1:4
y = (1/5)z. i.e y/z= 1/5 or y:z = 1:5
x = (1/4)y = (1/4)*(1/5)z = (1/20)z
x:z = 1:20
x:y:z = 1:4:20
Alternately, x:y = 1:4
y:z = 1:5 = 4:20 (multiplying both sides of the ratio by 5, which does not change the ratio)
so x:y:z = 1:4:20
:)